### v1

@@ Abstract
We investigate (3+1)d topological orders in fermionic systems with an anomalous $\mathbb{Z}_{2N}^{\rm F}$ symmetry, where its $\mathbb{Z}_{2}^{\rm F}$ subgroup is the fermion parity. Such an anomalous symmetry arises as the discrete subgroup of the chiral U(1) symmetry of $\nu$ copies of Weyl fermions of the same chirality. Guided by the crystalline correspondence principle, we construct (3+1)d symmetry-preserving gapped states on the boundary of a closely related (4+1)d $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ symmetry-protected topological (SPT) state (with $C_{N}$ being the $N$-fold rotation), whenever it is possible. In particular, for $\nu=N$, we show that the (3+1)d symmetric gapped state admits a topological $\mathbb{Z}_{4}$ gauge theory description at low energy, and propose that a similar theory saturates the corresponding $\mathbb{Z}_{2N}^{\rm F}$ anomaly. For $N\nmid\nu$, our construction cannot produce any topological quantum field theory (TQFT) symmetric gapped state; but for $\nu=N/2$, we find a non-TQFT symmetric gapped state via stacking lower-dimensional (2+1)d non-discrete-gauge-theory topological orders inhomogeneously. For other values of $\nu$, no symmetric gapped state is possible within our construction, which is consistent with the theorem by Cordova-Ohmori.
@@ I Introduction
’t Hooft anomaly is a fundamental property of global symmetry in quantum many-body systems and quantum field theories. It measures mild violations of gauge invariance in response to external gauge fields, and is an invariant under renormalization group flow (i.e. independent of energy scales). This property makes it a powerful tool to constrain the low-energy physics of an interacting system and has found many applications in both condensed matter and high-energy theories. Physically, ’t Hooft anomalies can be characterized by an invertible topological field theory (iTFT) in one dimension higher, i.e. a symmetry-protected topological (SPT) phase. This bulk-boundary correspondence has proven instrumental in our understanding of the classification of SPT phases.
In general, an anomalous system does not allow a completely featureless ground state. In dimensions higher than (1+1)d, there are three possible categories of low-energy dynamics:
@@
More recently, examples for non-perturbative global anomaly, often of finite group symmetries, have been systematic investigated. [^1] Perturbative local anomaly is detected by small (i.e. local) continuous symmetry transformation connected to the identity transformation, hence it is sometimes called the continuous anomaly.
Non-perturbative global anomaly is detected by large (i.e. global) discrete symmetry transformation disconnected from the identity transformation, hence it is sometimes called the discrete anomaly.
@@
Notably, this was shown to be the case for certain $\mathbb{Z}_{N}$ chiral anomaly in (3+1)d Córdova and Ohmori (2020), which can arise in Weyl fermions with chiral symmetries. Besides their fundamental roles in particle physics, Weyl fermions can also emerge in solid state systems, i.e. the so-called Weyl semimetals Wan et al. (2011). Alternatively, they can be viewed as boundary theories of (4+1)d fermionic SPT (FSPT) phases.
@@
In this work, we study (3+1)d gapped topological states with $\mathbb{Z}_{2N}^{\rm F}$ chiral anomaly. Suppose the anomaly class is indexed by $\nu$, it can be realized by $\nu$ copies of Weyl fermions $\psi_{\alpha}$ where $\alpha=1,2,\dots,\nu$, and the symmetry is generated by $\psi_{\alpha}\rightarrow e^{\hskip 1.0pt\mathrm{i}\hskip 1.0pt\pi/N}\psi_{\alpha}$ Córdova and Ohmori (2020). The simplest nontrivial example is $N=2$, where the anomaly is classified by the bordism group Freed and Hopkins (2016) $\Omega_{5}^{{\rm Spin}\times_{\mathbb{Z}_{2}^{\rm F}}{\mathbb{Z}_{4}}}=\mathbb{Z}_{16}$ Tachikawa and Yonekura (2018); Garcia-Etxebarria and Montero (2019); Hsieh (2018); Wan and Wang (2020); Guo et al. (2020), i.e. $\nu$ is defined mod 16.
@@
Interestingly, such a $\mathbb{Z}_{16}$ class shows up as the mixed gauge-gravity nonperturbative global anomaly Garcia-Etxebarria and Montero (2019); Wan and Wang (2020); Wang (2020a, b, 2021); Wang et al. (2022); Putrov and Wang (2023) for the Standard Model (SM) with a $\mathbb{Z}_{4,X}$ chiral symmetry, where $X$ is a particular combination of the baryon minus lepton $\mathbf{B}-\mathbf{L}$ number and the properly quantized electroweak hypercharge $\tilde{Y}$, i.e. $X\equiv 5(\mathbf{B}-\mathbf{L})-\frac{2}{3}\tilde{Y}$ proposed by Wilczek-Zee Krauss and Wilczek (1989a). The $\mathbb{Z}_{4,X}$ symmetry-preserving deformation can still violate baryon B conservation (which triggers nucleon decays) or lepton L conservation Weinberg (1979); Wilczek and Zee (1979). Note that $X^{2}=(-1)^{\rm F}$ and ${\rm U}(1)_{X}\supset\mathbb{Z}_{4,X}\supset\mathbb{Z}_{2}^{\rm F}$, thus $\mathbb{Z}_{4,X}=\mathbb{Z}_{4}^{\rm F}$. The three generations of experimentally confirmed quarks and leptons contribute a total number of Weyl fermions in the SM as $3\times 15=45$, which gives a $\mathbb{Z}_{4,X}$ anomaly index $\nu=45\mod 16=-3\mod 16$ (more details on this anomaly in the SM and the phenomenological implications can be found in Appendix B).
We will consider whether this kind of anomaly can be saturated by a symmetric gapped state without breaking the $\mathbb{Z}_{2N}^{\rm F}$ symmetry. Assuming that the symmetric gapped state is described by a TQFT, the constraint derived by Cordova and Ohmori in [Córdova and Ohmori, 2020] showed that this is not possible when $N\nmid\nu$. On the other hand, when $N\mid\nu$, no such obstructions are known to exist. It is an interesting and important question to understand what types of topological orders can saturate the anomaly. In addition, one would naturally like to know what is the “simplest” topological order satisfying the anomaly constraint.
@@
To address this question, we consider a UV deformation of the Weyl fermions by a spatially-inhomogeneous mass term. It breaks the $\mathbb{Z}_{2N}^{\rm F}$ symmetry, but preserves a $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ symmetry. Here $C_{N}$ denotes the $N$-fold rotation symmetry in (3+1)d. The resulting $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ symmetry is still anomalous, which manifests as chiral fermions appearing on the rotation axis (while the fermions are gapped away from the axis). The number of chiral fermions is precisely given by the anomaly index $\nu$. This is not a coincidence: these two symmetry groups, $\mathbb{Z}_{2N}^{\rm F}$ and $C_{N}\times\mathbb{Z}_{2}^{\rm F}$, are known to be related via the “crystalline correspondence principle” Thorngren and Else (2018); Cheng and Wang (2022); Zhang et al. (2020); Manjunath et al. (2023). Namely, the classifications of topological theories with these symmetries are always isomorphic. For SPT phases, the classification turns out to be $\mathbb{Z}_{16}$ for both cases when $N=2$.
To gap out the chiral fermions, we generalize the approach in [Yang and Cheng, 2024] to provide explicit constructions of gapped states for $\nu$ being a multiple of $N$. In particular, we find that $\nu=N$ anomaly can be saturated by a $\mathbb{Z}_{4}$ gauge theory. We then show from the explicit construction that the $C_{N}$ symmetry acts projectively on $\mathbb{Z}_{4}$ gauge charges, and the action on $\mathbb{Z}_{4}$ flux loops is essentially implemented by a static network of invertible topological defects. Via the crystalline correspondence principle, it suggests that the same TQFT realizes the $\mathbb{Z}_{2N}^{\rm F}$ anomaly when the symmetry is implemented by the same kind of invertible topological defects.
In addition, we demonstrate that $\mathbb{Z}_{2}$ gauge theory does not work in our construction, suggesting that the $\mathbb{Z}_{4}$ gauge theory is the “minimal” topological theory compatible with the anomaly.
For $\nu\neq 0\pmod{N}$, the Cordova-Ohmori constraint (applied to $\mathbb{Z}_{2N}^{\rm F}$) forbids any TQFT states. Interestingly, for $C_{N}\times\mathbb{Z}_{2}^{\rm F}$, we are able to construct a fully gapped ground state for $\nu=N/2$, but the state is highly anisotropic and cannot be “lifted” to a TQFT. For other values of $\nu$, we prove that our construction for $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ does not yield any fully gapped state, consistent with the $\mathbb{Z}_{2N}^{\rm F}$ result.
@@ I.1 $\mathbb{Z}_{2N}^{\rm F}$ symmetry and the nonperturbative global anomaly
The $\mathbb{Z}_{2N}^{\rm F}$ symmetry is still anomalous. Again by anomaly inflow, $\mathbb{Z}_{2N}^{\rm F}$ anomalies can be classified by $\mathbb{Z}_{2N}^{\rm F}$ fermionic SPT phases in (4+1)d. We are mostly interested in $N=2^{p}$. In this case, the Freed-Hopkins classification Freed and Hopkins (2016) of (4+1)d $\mathbb{Z}_{2N}^{\rm F}$ SPTs is given by the 5d bordism group of ${\rm Spin}^{\mathbb{Z}_{2N}}\equiv{\rm Spin}\times_{\mathbb{Z}_{2}^{\rm F}}{\mathbb{Z}_{2N}}\equiv\frac{{\rm Spin}\times{\mathbb{Z}_{2N}}}{\mathbb{Z}_{2}^{\rm F}}$ symmetry  Hsieh (2018); Guo et al. (2020):  [^2] Note that for a product of abelian groups, for clarity, we use $\times$ for the symmetry groups, while the direct sum $\oplus$ for the group classification of SPT states.
$\Omega_{5}^{{\rm Spin}^{\mathbb{Z}_{2N}}}=\mathbb{Z}_{2^{p+3}}\oplus\mathbb{Z}_{2^{p-1}},$
@@
which matches our calculation of the $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ FSPT classification using block state construction in Appendix A. For general $N$, we can always write $N=q\cdot 2^{p}$ where $q$ is an odd integer, and the symmetry group factorizes as $\mathbb{Z}_{2N}^{\rm F}=\mathbb{Z}_{2^{p+1}}^{\rm F}\times\mathbb{Z}_{q}$. In this case, the classification turns out to be
$\Omega_{5}^{{\rm Spin}^{\mathbb{Z}_{2^{p+1}}}}\oplus\Omega_{5}^{\rm Spin}(B\mathbb{Z}_{q}),$
@@
where the latter group is $\Omega_{5}^{\rm Spin}(B\mathbb{Z}_{q})\cong\Omega_{5}^{\rm SO}(B\mathbb{Z}_{q})$ isomorphic to the bosonic SPT phases protected by $\mathbb{Z}_{q}$ symmetry. TQFTs on the boundary of bosonic SPT phases, especially the “beyond-group-cohomology” ones, have been systematically studied in [Yang and Cheng, 2024]. In this work we focus on the $N=2^{p}$ case, where the fermionic part (from $\mathbb{Z}_{2^{p+1}}^{\rm F}$ symmetry) plays a role.
@@
However, not all of the (4+1)d $\mathbb{Z}_{2N}^{\rm F}$ SPT phases admit a gapped boundary TQFT description. To be concrete, denote the anomaly index by $\nu$, which can be realized by $\nu$ copies of Weyl fermions. In [Córdova and Ohmori, 2020], Cordova and Ohmori showed that when $N\nmid\nu$, a unitary spin TQFT cannot exist as the boundary theory to saturate the anomaly discussed above. We now briefly review their argument. Consider the SPT theory on $M_{4}\times\mathbb{R}_{+}$, where $M_{4}$ is a 4D spin manifold with Pontryagin number 48, and $\mathbb{R}_{+}$ is the upper half line +=$\mathbb{R}_{+}=\{x^{5}|x^{5}\geqslant 0\}$. Here the nontrivial SPT phase lives in the space $x^{5}>0$, while the trivial phase lives in $x^{5}<0$. The boundary theory $\mathcal{B}$ lives on $x^{5}=0$. Let $Z_{\mathcal{B}}[M_{4}]$ be the path integral of $\mathcal{B}$ on the interval $0\leqslant x^{5}\leqslant x$, $x>0$, which defines a boundary state $|\mathcal{\mathcal{B}}\rangle$ on $M_{4}\times\{x\}$. Their analysis shows that if $\mathcal{B}$ is a symmetry-preserving boundary theory, then when $N\nmid\nu$ we must have
@@
I.2 Crystalline correspondence
Starting from the (3+1)d Weyl fermion, we turn on an “s-wave” superconducting pairing, which completely gaps out the Weyl fermion. More explicitly, the superconducting Weyl fermion is described by the following Hamiltonian:
[…]

### v2

@@ Abstract
We investigate (3+1)d topological orders in fermionic systems with an anomalous $\mathbb{Z}_{2N}^{\mathrm{F}}$ symmetry, where its $\mathbb{Z}_{2}^{\mathrm{F}}$ subgroup is the fermion parity. Such an anomalous symmetry arises as the discrete subgroup of the chiral U(1) symmetry of $\nu$ copies of Weyl fermions of the same chirality. Inspired by the crystalline correspondence principle, we deform the anomalous $\mathbb{Z}_{2N}^{\mathrm{F}}$ symmetry of (3+1)d Weyl fermion to the anomalous $C_{N}\times\mathbb{Z}_{2}^{\mathrm{F}}$ symmetry. Then we microscopically construct symmetry-preserving gapped boundary states of the closely-related (4+1)d $C_{N}\times\mathbb{Z}_{2}^{\mathrm{F}}$ symmetry-protected topological (SPT) state (with $C_{N}$ being the $N$-fold rotation), whenever it is possible. In particular, for $\nu=N$, we show that the (3+1)d symmetric gapped state admits a topological $\mathbb{Z}_{4}$ gauge theory description at low energy, and propose that a similar theory saturates the corresponding $\mathbb{Z}_{2N}^{\mathrm{F}}$ anomaly. For $N\nmid\nu$, our construction admits no topological quantum field theory (TQFT) symmetric gapped state; while for $\nu=N/2$, we find a non-TQFT symmetric gapped state via stacking lower-dimensional (2+1)d non-discrete-gauge-theory topological order inhomogeneously. For other values of $\nu$, no symmetric gapped state is possible within our construction, which is consistent with the no-go theorem by Cordova-Ohmori.
@@ I Introduction
’t Hooft anomaly is a fundamental property of global symmetry in quantum many-body systems and quantum field theories. It measures mild violations of gauge invariance in response to external gauge fields, and is an invariant under renormalization group flow (i.e. independent of energy scales). This property makes it a powerful tool to constrain the low-energy physics of an interacting system and has found many applications in both condensed matter and high-energy theories. Physically, ’t Hooft anomalies can be characterized by an invertible topological field theory (iTFT) in one dimension higher, i.e. a symmetry-protected topological (SPT) phase. This bulk-boundary correspondence has proven instrumental in our understanding of the classification of SPTs.
In general, an anomalous system does not allow a completely featureless ground state. In spacetime dimensions higher than (1+1)d, there are three possible categories of low-energy dynamics:
@@
More recently, examples for non-perturbative global anomaly, often of finite group symmetries, have been systematically investigated. [^1] Perturbative local anomaly is detected by small (i.e. local) continuous symmetry transformation connected to the identity transformation, hence it is sometimes called the continuous anomaly. Non-perturbative global anomaly is detected by large (i.e. global) discrete symmetry transformation disconnected from the identity transformation, hence it is sometimes called the discrete anomaly. Notably, this was shown to be the case for certain $\mathbb{Z}_{N}$ chiral anomaly in (3+1)d Córdova and Ohmori (2020), which can arise in Weyl fermions with chiral symmetries. Besides their fundamental roles in particle physics, Weyl fermions can also emerge in solid state systems, i.e. the so-called Weyl semimetals Wan et al. (2011). Alternatively, they can be viewed as boundary theories of (4+1)d fermionic SPT (FSPT) phases.
@@
In this work, we study (3+1)d gapped topological states with $\mathbb{Z}_{2N}^{\rm F}$ chiral anomaly as the boundary theory of (4+1)d $\mathbb{Z}_{2N}^{\rm F}$ FSPT. Suppose the anomaly class is indexed by $\nu$, it can be realized by $\nu$ copies of Weyl fermions $\psi_{\alpha}$ where $\alpha=1,2,\dots,\nu$, and the symmetry is generated by $\psi_{\alpha}\rightarrow e^{\hskip 1.0pt\mathrm{i}\hskip 1.0pt\pi/N}\psi_{\alpha}$ Córdova and Ohmori (2020). The simplest nontrivial example is $N=2$, where the anomaly is classified by the bordism group Freed and Hopkins (2016) $\Omega_{5}^{{\rm Spin}\times_{\mathbb{Z}_{2}^{\rm F}}{\mathbb{Z}_{4}}}=\mathbb{Z}_{16}$ Tachikawa and Yonekura (2018); García-Etxebarria and Montero (2019); Hsieh (2018); Wan and Wang (2020); Guo et al. (2020), i.e. $\nu$ is defined mod 16.
@@
Interestingly, such a $\mathbb{Z}_{16}$ class shows up as the mixed gauge-gravity non-perturbative global anomaly García-Etxebarria and Montero (2019); Wan and Wang (2020); Wang (2020a, b, 2021); Wang et al. (2022a, b); Putrov and Wang (2024) for the Standard Model (SM) with a $\mathbb{Z}_{4,X}$ chiral symmetry, where $X$ is a particular combination of the baryon minus lepton $\mathbf{B}-\mathbf{L}$ number and the properly quantized electroweak hypercharge $\tilde{Y}$, i.e. $X\equiv 5(\mathbf{B}-\mathbf{L})-\frac{2}{3}\tilde{Y}$ proposed by Wilczek-Zee Krauss and Wilczek (1989a). The $\mathbb{Z}_{4,X}$ symmetry-preserving deformation can still violate baryon B conservation (which triggers nucleon decays) or lepton L conservation Weinberg (1979); Wilczek and Zee (1979). Note that $X^{2}=(-1)^{\rm F}$ and ${\rm U}(1)_{X}\supset\mathbb{Z}_{4,X}\supset\mathbb{Z}_{2}^{\rm F}$, thus $\mathbb{Z}_{4,X}=\mathbb{Z}_{4}^{\rm F}$. The three generations of experimentally confirmed quarks and leptons contribute a total number of Weyl fermions in the SM as $3\times 15=45$, which gives a $\mathbb{Z}_{4,X}$ anomaly index $\nu=45\mod 16=-3\mod 16$ (more details on this anomaly in the SM and the phenomenological implications can be found in Appendix B).
We then address the following question:
What is the minimal topological order saturating the $\mathbb{Z}_{2N}^{\rm F}$ anomaly?
Microscopically, we consider whether the $\mathbb{Z}_{2N}^{\rm F}$ anomaly can be saturated by a symmetric gapped state as the boundary theory of (4+1)d $\mathbb{Z}_{2N}^{\rm F}$ FSPT, realized via the block state construction Song et al. (2017). Assuming that the symmetric gapped state at the low energy is described by a TQFT, the no-go theorem shown by Cordova and Ohmori in [Córdova and Ohmori, 2020] suggests that when $N\nmid\nu$ there exists no gapped boundary TQFT description. On the other hand, when $N\mid\nu$, no such obstructions should exist. Thus we would like to understand what topological orders are sufficient to saturate the anomaly. In particular, we provide a microscopic construction of the symmetric gapped boundary state, which can be unfolded in the following steps:
  1. 1.
@@
    We start from (3+1)d Weyl fermions. They have anomalous $\mathbb{Z}_{2N}^{\rm F}$ symmetry and can be viewed as the boundary states of a (4+1)d $\mathbb{Z}_{2N}^{{\rm F}}$ FSPT. In Section I.2, we consider a UV deformation of the Weyl fermions by a spatially-inhomogeneous mass term. It breaks the $\mathbb{Z}_{2N}^{\rm F}$ symmetry while preserving a $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ symmetry. The resulting $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ symmetry is still anomalous, which manifests as chiral fermions appearing on the (1+1)d rotational axis (while the fermions are gapped away from the axis). The number of chiral fermions signaling the anomalous $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ symmetry is precisely given by the anomaly index $\nu$ of $\mathbb{Z}_{2N}^{\rm F}$ symmetry of the Weyl fermion. In fact, these two symmetry groups, $\mathbb{Z}_{2N}^{\rm F}$ and $C_{N}\times\mathbb{Z}_{2}^{\rm F}$, are known to be related via the “crystalline correspondence principle” Thorngren and Else (2018); Cheng and Wang (2022); Zhang et al. (2020); Manjunath et al. (2023). Namely, the classifications of invertible TQFTs with these symmetries are isomorphic. For SPT phases, the classification turns out to be $\mathbb{Z}_{16}$ for both cases when $N=2$.
  2. 2.
[…]
    $1\rightarrow G_{A}\rightarrow C_{N^{\prime}}\rightarrow C_{N}\rightarrow 1\,.$
[…]
    Within our construction, we give a slight refinement of Cordova-Ohmori’s no-go theorem:
    • for $\nu=N$, i.e. FSPTs characterized by Majorana chain decoration, the minimal topological order saturating the $\mathbb{Z}_{2N}^{\rm F}$ anomaly is a $\mathbb{Z}_{4}$ gauge theory;
    • for $\nu=\frac{N}{2}$, i.e. FSPTs characterized by $p_{x}+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp_{y}$ decoration, there exists no gapped boundary TQFT description but a highly-anisotropic symmetric gapped boundary state. We conjecture that same is true for boundary states of other (4+1)d FSPTs with $p_{x}+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp_{y}$ decoration;
    • for other values of $N$ our construction shows no symmetric gapped boundary states, consistent with the no-go theorem, and we conjecture that the boundary state must be gapless or $C_{N}$ symmetry breaking.
@@ I.1 $\mathbb{Z}_{2N}^{\rm F}$ symmetry and the nonperturbative global anomaly
The $\mathbb{Z}_{2N}^{\rm F}$ symmetry is still anomalous. Again by anomaly inflow, $\mathbb{Z}_{2N}^{\rm F}$ anomalies can be classified by $\mathbb{Z}_{2N}^{\rm F}$ fermionic SPT phases in (4+1)d. We are mostly interested in $N=2^{p}$. In this case, the Freed-Hopkins classification Freed and Hopkins (2016) of (4+1)d $\mathbb{Z}_{2N}^{\rm F}$ SPTs is given by the 5d bordism group $\Omega_{5}$ of
${\rm Spin}^{\mathbb{Z}_{2N}}\equiv{\rm Spin}\times_{\mathbb{Z}_{2}^{\rm F}}{\mathbb{Z}_{2N}}\equiv\frac{{\rm Spin}\times{\mathbb{Z}_{2N}}}{\mathbb{Z}_{2}^{\rm F}}$
symmetry  Hsieh (2018); Guo et al. (2020):  [^2] Note that for a product of abelian groups, for clarity, we use $\times$ for the symmetry groups, while the direct sum $\oplus$ for the group classification of SPT states.
$\Omega_{5}^{{\rm Spin}^{\mathbb{Z}_{2^{p+1}}}}=\mathbb{Z}_{2^{p+3}}\oplus\mathbb{Z}_{2^{p-1}},$
@@
which matches our calculation of the $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ FSPT classification using block state construction in Appendix A. For general $N$, we can always write $N=2^{p}\cdot q$ where $q$ is a generic odd integer, and the symmetry group factorizes as $\mathbb{Z}_{2N}^{\rm F}=\mathbb{Z}_{2^{p+1}}^{\rm F}\times\mathbb{Z}_{q}$. In this case, the classification turns out to be
$\Omega_{5}^{{\rm Spin}^{\mathbb{Z}_{2^{p+1}\cdot q}}}$ $=$ $\Omega_{5}^{{\rm Spin}^{\mathbb{Z}_{2^{p+1}}}}\oplus\Omega_{5}^{\rm Spin}(B\mathbb{Z}_{q})$ (1)
$\cong$ $\Omega_{5}^{{\rm Spin}^{\mathbb{Z}_{2^{p+1}}}}\oplus\tilde{\Omega}_{5}^{\rm SO}(B\mathbb{Z}_{q})$ (2)
where the latter group is $\Omega_{5}^{\rm Spin}(B\mathbb{Z}_{q})\cong\tilde{\Omega}_{5}^{\rm SO}(B\mathbb{Z}_{q})$ isomorphic to the bosonic SPT phases protected by $\mathbb{Z}_{q}$ symmetry. Here $\tilde{\Omega}_{5}^{\rm SO}(\mathrm{B}G)\equiv\Omega_{5}^{\rm SO}(\mathrm{B}G)/\Omega_{5}^{\rm SO}$ is the reduced bordism group, modding out the $\Omega_{5}^{\rm SO}=\Omega_{5}^{\rm SO}(pt)$. Moreover, if we write $N=2^{p}\cdot q=2^{p}\cdot 3^{r}\cdot s$ where $s$ is a generic odd integer that contains no factor of 3, then
$\Omega_{5}^{{\rm Spin}^{\mathbb{Z}_{2^{p+1}\cdot 3^{r}\cdot s}}}\cong\Omega_{5}^{{\rm Spin}^{\mathbb{Z}_{2^{p+1}}}}\oplus\tilde{\Omega}_{5}^{\rm SO}(B\mathbb{Z}_{3^{r}\cdot s})$ (3)
$=\mathbb{Z}_{2^{p+3}}\oplus\mathbb{Z}_{2^{p-1}}\oplus\mathbb{Z}_{3^{r+1}}\oplus\mathbb{Z}_{3^{r-1}}\oplus\mathbb{Z}_{s}\oplus\mathbb{Z}_{s}.{}$
@@
TQFTs on the boundary of bosonic SPT phases, especially the “beyond-group-cohomology” ones, have been systematically studied in [Yang and Cheng, 2024]. Hence, in this work, we focus on the $N=2^{p}$ case, where the fermionic part (from $\mathbb{Z}_{2^{p+1}}^{\rm F}$ symmetry) plays a role.
@@
However, not all of the (4+1)d $\mathbb{Z}_{2N}^{\rm F}$ SPT phases admit a gapped boundary TQFT description. To be concrete, denote the anomaly index by $\nu$, which can be realized by $\nu$ copies of Weyl fermions. In [Córdova and Ohmori, 2020], Cordova and Ohmori showed that when $N\nmid\nu$, a unitary spin TQFT cannot exist as the boundary theory to saturate the anomaly discussed above. We now briefly review their argument. Consider the SPT theory on $M_{4}\times\mathbb{R}_{+}$, where $M_{4}$ is a 4D spin manifold [^3] We denote nnD for space dimension, and $(n+1)d$ for spacetime dimension. with Pontryagin number 48, and $\mathbb{R}_{+}$ is the upper half line $\mathbb{R}_{+}=\{x^{5}|x^{5}\geqslant 0\}$. Here the nontrivial SPT phase lives in the space $x^{5}>0$, while the trivial phase lives in $x^{5}<0$. The boundary theory $\mathcal{B}$ lives on $x^{5}=0$. Let $Z_{\mathcal{B}}[M_{4}]$ be the path integral of $\mathcal{B}$ on the interval $0\leqslant x^{5}\leqslant x$, $x>0$, which defines a boundary state $|\mathcal{\mathcal{B}}\rangle$ on $M_{4}\times\{x\}$. Their analysis shows that if $\mathcal{B}$ is a symmetry-preserving boundary theory, then when $N\nmid\nu$ we must have
@@
I.2 $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ symmetry
In this section, we discuss the UV deformation from the anomalous $\mathbb{Z}_{2N}^{\rm F}$ symmetry to the anomalous $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ symmetry. Starting from the (3+1)d Weyl fermion, we turn on an “s-wave” superconducting pairing, which completely gaps out the Weyl fermion. More explicitly, the superconducting Weyl fermion is described by the following Hamiltonian:
$H=\int\hskip 1.0pt\mathrm{d}^{3}\mathbf{x}\left[\psi^{\dagger}(\mathbf{x})\bm{\sigma}\cdot\mathbf{p}\psi(\mathbf{x})+\Delta_{0}\psi^{\dagger}(\mathbf{x})\hskip 1.0pt\mathrm{i}\hskip 1.0pt\sigma_{y}\psi^{\dagger}(\mathbf{x})+\text{h.c.}\right].$